Applications of Electronics of Differential Equation in Physics


1
Expert's answer
2022-03-25T10:26:30-0400

Differential equations are useful in physics as the help solve alot of electronic problems.

Examples include;

  1. Ohm's Law; "V=IR"
  2. Voltage across an Inductor; "V=L\frac{dI}{dt}"
  3. Current through a capacitor; "I=C\frac{dV}{dt}"
  4. Gauss Law; "\Delta.D^\rightarrow=\rho_v"
  5. Ampere's circuit law; "\Delta.\Eta^\rightarrow=J^\rightarrow+\frac{dD^\rightarrow}{dt}"